ANALYSIS OF PERTURBATIONS OBSERVED IN THE FT SPECTRA OF CuCl_{2}

\underline{A. YIANNOPOULOU}, A. J. ROSS, P. CROZET, Laboratoire de Spectrométrie Ionique et Moléculaire, CNRS et Université Lyon-I (UMR 5579), Campus la Doua, 69622 Villeurbanne Cedex, France; and J. M. BROWN, Physical and Theoretical Chemistry Laboratory, Oxford University, South Parks Road, Oxford, OX1 3QZ, England.

Abnormally large e-f separations have been observed in the (v_{1}, 0, v_{3}) levels of the ground X^{2}Π_{g(3/2)} state of CuCl_{2} about 2500 cm^{−1} above the lowest vibrational level (0,0,0). The largest splittings are accompanied by extra lines in the dispersed fluorescence spectra. Some of the extra lines have resolvable Cu hyperfine structure; they are assigned as transitions to levels which result from mixing essentially between ^{2}Π_{g(1/2)} and a nearby ^{2}Σ_{g} state. We have used an effective Hamiltonian to model the^{2}Π ∼ ^{2}Σ interaction and derived molecular constants for the unknown 2Σ state using a non-linear least-squares fitting routine. Taking an arbitrary value of the spin-orbit parameter A_{Π} of -300 cm^{−1}, we estimate the ^{2}Σ state to lie around 3000 cm^{−1} above the (0,0,0) level of the X^{2}Π_{g(3/2)} ground state.